OpenStudy (anonymous):

2.What is the sum of the geometric sequence 4, 12, 36 … if there are 9 terms? (1 point)

OpenStudy (anonymous):

well do you know what the sequence is?

hartnn (hartnn):

and can you find out the common ratio (r) first ??

OpenStudy (anonymous):

because if you look 4*3 =12 12*3=36 so the sequence must be x*3=y so your sequence chain would be 4, 12, 36, 108, 324, 972, 2916, 8748, 26244. 4+12+36+108+324+972+2916+8748+26244= 39364 This is the answer as well as the steps to get said answer

hartnn (hartnn):

or you could use the formula for the sum of terms of a geometric sequence. This formula especially is useful when you have many terms \(\Large S_n = a_1 \dfrac{r^{n}-1}{r-1}\)

hartnn (hartnn):

a1 = first term = 4 here r= common ratio =...? n = number of terms = 9 here

OpenStudy (anonymous):

thanks @hartnn i had completely forgotten about that formula.

hartnn (hartnn):

oh, you're welcome ^_^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so it would be \[4(\frac{ 3^9-1 }{ 3-1 })\] = 39364

OpenStudy (anonymous):

thats a lot easier then the method i used